# MITO: Enabling Non-Line-of-Sight Perception using Millimeter-waves through Real-World Datasets and Simulation Tools

### v1

@@ Abstract
We present MITO, the first dataset of multi-spectral millimeter-wave (mmWave) images of everyday objects. Unlike visible light, mmWave signals can image through everyday occlusions (e.g., cardboard boxes, fabric, plastic). However, due to the dearth of publicly-available mmWave images and the interdisciplinary challenges in collecting and processing mmWave signals, it remains difficult today for computer vision researchers to develop mmWave-based non-line-of-sight perception algorithms and models.
To overcome these challenges, we introduce a real-world dataset and open-source simulation tool for mmWave imaging. The dataset is acquired using a UR5 robotic arm with two mmWave radars operating at different frequencies and an RGB-D camera. Through a radar signal processing pipeline, we capture and create over 580 real-world 3D mmWave images taken from over 76 different objects in the YCB dataset, a standard benchmarking dataset for robotics manipulation. For each object, we provide real-world mmWave images in line-of-sight and non-line-of-sight, as well as RGB-D images and ground truth segmentation masks. We also develop an open-source simulation tool that can be used to generate synthetic mmWave images for any 3D triangle mesh, which achieves a median F-Score of 94% when compared to real-world mmWave images.
We demonstrate the usefulness of this dataset and simulation tool in multiple computer vision tasks in non-line-of-sight. First, we perform object segmentation for mmWave images leveraging the segment anything model (SAM), and achieve a median precision and recall of 92.6% and 64%. Second, we train a classifier that can recognize objects in non-line-of-sight. It is trained on synthetic images and can classify in real-world scenarios with 85% accuracy.
The dataset of images, raw data, and the simulation – along with python and cuda scripts for downloading the dataset, processing the raw data, and visualizing the images – is all available open-source. We believe that MITO will serve as an instrumental resource that enables computer vision researchers to develop algorithms and models for non-line-of-sight perception using mmWaves, and hope that this first-of-its-kind dataset will have a similar impact as early camera-based datasets that paved the way for significant innovations in the field.
@@ 1 Introduction
However, relying primarily on optical perception (e.g., cameras, LiDARs, etc) inherently limits existing systems to line-of-sight. [^1] While some visible-light systems can rely on reflected laser light to image objects around the corner, they cannot image fully occluded items (e.g., within a closed box). For example, a camera-based system cannot see inside a closed box to confirm e-commerce orders are correct or detect broken objects during shipping. Nor is it feasible for camera-based robots to efficiently plan complex, multi-step tasks in practical environments where required objects may be fully occluded.
@@
Inspired by how early imaging datasets, such as MSRC[21], Pascal[20], Berkeley 3D[27], and YCB[17], paved the way for many advancements within computer vision, we believe that a dataset of mmWave images of everyday objects is fundamental for enabling pervasive NLOS perception.
In this paper, we present MITO.333mmWave Imaging Through Occlusions To the best of our knowledge, this is the first dataset of multi-spectral mmWave images for everyday objects. It builds on the YCB dataset, a standard dataset for robotic manipulation which already provides 2D and 3D camera scans of many common objects. In building MITO, we introduce a few key contributions:
@@
  • Real World Dataset. Using a custom-built robotic imaging system and signal processing pipeline, we capture images of 76 objects in the YCB dataset (out of a total of 80 objects). We captured over 580 real-world images of objects in both line-of-sight and non-line-of-sight. For each object, we capture mmWave images at multiple different frequencies. The frequency of the signal introduces an important tradeoff between the image resolution and the ability for the radars to image objects under thick occlusions. We also include the RGB-D images of the objects (when in LOS) and a ground truth object segmentation produced with the segment-anything model.
@@
  • Open-Source mmWave Simulation Tool. We also created an open-source simulation tool that can produce synthetic mmWave images for any 3D triangle mesh. Given the million of meshes available online[10, 1], this provides the flexibility to produce a significant synthetic dataset. Importantly, different materials interact differently with mmWave signals and therefore produce different images. Therefore, we build two different modeling methods into our simulation, the combination of which can be used to simulate a wide range of material properties. When compared to real-world mmWave images, our simulation achieves a median 3D F-Score of 94%.
@@
  • Example Computer Vision Tasks. Using MITO, we demonstrated two potential computer vision tasks in non-line-of-sight. First, we show object segmentation, using the segment-anything model (SAM)[28] and a mmWave power-based prompter. Second, we show object classification, which is trained entirely on synthetic images and is able to classify objects in real-world images with 85% accuracy.
The dataset, all code, and a video is available here:
https://github.com/signalkinetics/MITO_Codebase
@@ 2 Background
In this section, we describe how mmWave imaging works at a high level, and refer readers to [35] for more details. A mmWave radar transmits wireless signals with millimeter-wave wavelengths, captures their reflections, and uses them to generate 3D images, as shown in Fig. 3a. It uses two core signal processing techniques. First, it estimates distance using chirp signals. Specifically, by comparing the frequency between the transmitted and received signals it can estimate the distance of each reflection. Second, it measures the angle of a reflection (in both azimuth and elevation) by comparing the received signals between different antennas. By measuring the reflection power emanating from each voxel in 3D space (defined by distance, elevation angle, and azimuth angle), it can perform 3D reconstruction.
@@
where $I(x,y,z)$ is the complex mmWave image at location $(x,y,z)$, $K$ is the number of antennas, $N$ is the number of samples in each chirp, $S_{j,k}$ is the jth sample of the chirp from the kth antenna, and $\lambda_{j}$ is the wavelength of the signal at sample j, which is proportional to the speed of light c and the frequency f as $\lambda=\frac{c}{f}$. $d(x,y,z,k)$ is the round-trip distance from the kth antenna (at location $p_{k}$) to the point $(x,y,z)$ and is defined as $d(x,y,z,k)=2||(x,y,z)-p_{k}||$.
2.1.1 Resolution.
One important characteristic of a mmWave image is its resolution. The resolution in range (depth) is determined by the bandwidth, or the range of frequencies that the chirp covers. We compute the depth resolution $\delta_{z}$ as:
@@
2.1.2 Synthetic Aperture Radar (SAR) Imaging.
@@
2.2 Impact of Frequency.
One important question is what frequency to use for imaging. The frequency introduces an important tradeoff. First, higher frequencies result in higher resolution (as seen in Eq. 3). On the other hand, higher frequency signals experience more attenuation, or power loss, as they travel through materials[8]. This means that high frequency signals may not be able to image through as thick of occlusions compared to lower frequencies. This is why MITO’s dataset and simulation tool include multi-spectral mmWave images (i.e., at different frequencies).
2.3 Types of Reflections.
mmWave signals primarily exhibit specular (mirror-like) reflections[33]. This characteristic is illustrated in Fig. 3. Due to the significantly longer wavelengths of mmWave signals compared to visible light (millimeters vs nanometers), surfaces that appear diffuse (i.e., omnidirectional) for visible light (blue) often appear specular for mmWave signals (red).
In addition to specular reflections, mmWaves also exhibit strong reflections when a wave strikes a sharp edge of an object, scattering it in multiple directions. This type of reflection, known as edge diffraction, can impact signal propagation and perception differently from diffuse or specular reflections.
@@
3 MITO: Properties and Structure
In this section, we describe the contents of MITO.
Objects. MITO builds on the YCB dataset[17], a standard dataset of objects used for robotic manipulation. The YCB dataset already provides 2D images and 3D scans of 80 everyday objects that we use to simulate synthetic mmWave images of the objects. We have included 76 of these objects in MITO [^4] The remaining 4 items were discarded mainly because we could not obtain the exact items from the YCB dataset to image them..
Object Diversity. Since our dataset represents the vast majority of the YCB dataset, our objects are diverse across shape, size, material, and type. In terms of shape, our dataset includes spherical, flat, curved, and complex shapes. In terms of size, it includes objects whose dimensions range from 1.5cm to 47cm. Our objects include various materials such as wood, plastic (e.g., HDPE, PETE, Polypropelene, Nylon, etc), metal, cardboard, rubber, as well as objects with multiple materials. Additionally, our objects represent a wide range of categories, including household tools (wrench, screwdriver, hammer), groceries (sugar box, soup can, cracker box), kitchen supplies (spatula, plate, skillet), toys (marbles, rubik’s cube, baseball), and office supplies (scissors, dry erase markers) .
Real-World Multi-Spectral mmWave Images. We provide over 580 3D, complex-valued mmWave images in both line-of-sight (LOS) and non-line-of-sight (NLOS) settings, and at two different frequencies (24GHz and 77GHz).
Recall from Sec. 2.2 that the frequency of the mmWave signal impacts both the resolution of the resulting image and the ability for the signal to traverse through occlusions. For example, we show four example images in Fig. 4. Fig.4a-b show a 77GHz and 24GHz image of a padlock in LOS, respectively. The 77GHz image is higher resolution, as expected. On the other hand, Fig. 4c-d show 77GHz and 24GHz images of the same object in NLOS (placed underneath cardboard). In this case, the 24GHz image is very similar to the LOS case, while the 77GHz shows some artifacts that were not present in the LOS case. This is because the 24GHz image travels through the cardboard relatively unaffected, while the 77GHz is partially reflected and absorbed by the cardboard. It is our hope that by including two different frequencies, future algorithms can leverage both images to benefit from both high-resolution and robustness to thick occlusions.
Raw mmWave Data. In addition to the processed mm-Wave images, we also include the raw signals collected during each experiment, as well as the radar’s location for each measurement. The code for processing raw measurements into a mmWave image is included as well. This allows researchers to investigate many different avenues, such as experimenting with other imaging algorithms or evaluating the impact of removing measurements from an experiment.
RGB-D Images. For each mmWave image, we also take a corresponding RGB-D image of the object. In the case of a NLOS image, we take the camera image before occluding the object. We also provide the ground truth segmentation mask using the RGB-D images. The camera’s location is also recorded, which allows us to transform the segmentation masks to the same frame as the mmWave images. These transformed masks are also included and can be used to evaluate segmentation and object completion networks.
Synthetic mmWave Images. In addition to all the real-world data described above, we also provide a simulation tool to generate synthetic mmWave images, which can be used to increase the volume of data to aid in model training. We include at least 2 different simulation images for each object, along with the simulation tool for producing additional images (See Sec. 5 for more details).
4 Data Collection and Validation
Next, we describe how we collect the data in MITO.
4.1 Real-World Data Collection
Object Placement. Similar to how initial RGB datasets provided images of objects with white backgrounds, our goal in the initial images of MITO is to create a simple image of just the object without a complex background. To do so, we image the objects in two setups, each of which minimizes background reflections in a different way. In the first, we place each object on a tripod [^5] We hot-glue a 3D printed nut to each object, and screw it onto a threaded tripod. such that there is no surface directly below the object, preventing any strong reflections from the background. In the second, we place the object directly on a styrofoam block, since styrofoam is largely transparent to mmWave signals[11].
Data Collection Setup. We build a robotic system capable of mmWave imaging. We connect a 77GHz radar (TI IWR1443Boost, $<$\$300[9] and DCA1000EVM, $<$\$600[2]) and a 24GHz radar (Infineon Position2Go, $<$\$300[7]), and Intel Realsense D415[26] to an end-effector of a UR5e robotic arm[46] using custom-designed 3D printed parts. At the start of an experiment, the robot moves the camera to the middle of the workspace to take an RGB-D image. Then, it moves the radars throughout the workspace, covering a rectangular space. The aperture of our experiments range from 55cm by 22cm to 55cm by 29cm. [^6] Each experiment uses an aperture which covers the full object. The radars continuously measure mmWave signals while the robot is moving, and the robot records its trajectory over time. After we complete an experiment in LOS, we run an experiment in NLOS without moving the object. We place a layer of cardboard over the object to simulate the case where the object was placed inside a box.
Robot - Radar Synchronization. To process the mmWave image, we need to know the location that each measurement was taken (per Eq. 1). To do this, we interpolate the timestamped robot locations to match the timestamp of the radar measurements. Since the robot and radars are controlled from different computers, we set both computers to the same NTP time server to ensure accurate time synchronization.
mmWave Image Processing. Next, the raw measurements are processed into a 3D mmWave image by applying Eq. 1. Example images can be seen in Fig. 5, which shows the segmented RGB images (1st column), 77GHz LOS (2nd column), 77GHz NLOS (3rd column), 24GHz LOS (4th column), and 24GHz NLOS (5th column) mmWave images. Additional images are provided in the supplementary.
Background Measurements. We also include experiments when the tripod/styrofoam is placed in the scene without any object. This can be used to remove the impact of the tripod/styrofoam from images by subtracting the empty image from the main image[13].
Ground Truth Segmentations. Given the RGB-D images of the objects in LOS, we produce ground-truth segmentation masks of the objects in each experiment using the segment-anything model (SAM). We use an interactive interface for SAM to select the points on the RGB picture to use as a prompt for the model. The human annotator continues selecting points until the mask covers the full object. We then align the RGB ground-truth mask with the mmWave image. The details for this alignment can be found in the supplementary material.
5 Simulating mmWave Images

### v3

# MITO: A Millimeter-Wave Dataset and Simulator for Non-Line-of-Sight Perception

@@ Abstract
The ability to observe the world is fundamental to reasoning and making informed decisions on how to interact with the environment. However, optical perception can often be disrupted due to common occurrences, such as occlusions, which can pose challenges to existing vision systems. We present MITO, the first millimeter-wave (mmWave) dataset of diverse, everyday objects, collected using a UR5 robotic arm with two mmWave radars operating at different frequencies and an RGB-D camera. Unlike visible light, mmWave signals can penetrate common occlusions (e.g., cardboard boxes, fabric, plastic) but each mmWave frame has much lower resolution than typical cameras. To capture higher-resolution mmWave images, we leverage the robot’s mobility and fuse frames over the synthesized aperture. MITO captures over 24 million mmWave frames and uses them to generate 550 high-resolution mmWave (synthetic aperture) images in line-of-sight and non-light-of-sight (NLOS), as well as RGB-D images, segmentation masks, and raw mmWave signals, taken from 76 different objects. We develop an open-source simulation tool that can be used to generate synthetic mmWave images for any 3D triangle mesh. Finally, we demonstrate the utility of our dataset and simulator for enabling broader NLOS perception by developing benchmarks for NLOS segmentation and classification.
@@ 1 Introduction
However, relying primarily on optical perception (e.g., cameras, LiDARs, etc) inherently limits existing systems to line-of-sight (LOS). [^1] While some visible-light systems can rely on reflected laser light to image objects around the corner, they cannot image fully occluded items (e.g., within a closed box). For example, a camera-based system cannot see inside a closed box to confirm e-commerce orders are correct or detect broken objects during shipping. Nor is it feasible for camera-based robots to efficiently plan complex, multi-step tasks in practical environments where required objects may be fully occluded.
@@
Inspired by how early imaging datasets, such as MSRC [21], Pascal [19], Berkeley 3D [26], and YCB [14], paved the way for many advancements within computer vision, we believe that a dataset of mmWave images of everyday objects is fundamental for enabling pervasive NLOS perception. To summarize, we make the following key contributions:
@@
  • Dataset. We create MITO, a novel dataset of mmWave images for everyday objects, using a custom-built robotic imaging system and signal processing pipeline. We capture over 24 million mmWave frames, and use them to generate 550 images of 76 objects from the YCB dataset (out of a total of 80 objects) in both line-of-sight and non-line-of-sight. For each object, we capture mmWave images at multiple different frequencies. We also provide RGB-D images of the objects (when in LOS) and ground truth object segmentations.
@@
  • Simulator. We develop an open-source simulator that can produce synthetic mmWave images for any 3D triangle mesh. Since different materials interact differently with mmWave signals and, therefore, produce different images, we build two different modeling methods into our simulation, the combination of which can be used to simulate a wide range of material properties.
@@
  • Non-line-of-sight Applications. Using MITO, we establish baselines for two computer vision tasks in non-line-of-sight. First, we show object segmentation, using the segment-anything model (SAM) [27] and a mmWave power-based prompter. Second, we show shape classification, which is trained entirely on synthetic images and can classify real-world images.
The dataset, code, and a video demonstration are available here:
2 Related Work
mmWave Datasets. Recognizing the importance of millimeter-wave perception, past research has made initial steps towards millimeter-wave datasets. However, the vast majority of them are focused either on imaging humans [15, 16, 46, 16, 37, 48] for action recognition, outdoor imaging for autonomous driving [11, 13, 22, 28, 34, 51, 35, 49, 45, 23], or imaging of building interiors (walls, hallways) [12, 36]. In the context of higher resolution imaging of objects, the only dataset that we are aware of is focused on guns and knives for TSA and detecting concealed weapons through body scanners for airport security applications [31, 39, 38]. Our work, however, is focused on generating a dataset of a larger, diverse set of everyday objects for application such as robotic manipulation and perception.
mmWave Image Simulators. Given the growth in millimeter wave devices for networking and sensing, the vast majority of existing millimeter wave simulations have also focused on understanding channel characteristics (e.g., for 5G networks) [8, 33, 41] or for imaging humans for action recognition and HCI applications [47, 52, 50]. In principle, these simulators could be used also for generating mmWave images of everyday objects; however, existing simulators are extremely expensive, costing tens of thousands of dollars [8] and/or are limited to simulating metallic items and cannot simulate everyday objects with different properties [38]. In contrast, we developed an open-source simulator and demonstrated that it generates synthetic data that matches real-world collected data. We believe that this tool combined with advances in vision algorithms can lead to very high accuracy in non-line-of-sight scenarios.
mmWave Perception. Prior work has demonstrated the use of millimeter waves for various perception tasks, including detection and classification. However, these approaches are limited to specific domains, such as vehicles [29, 22, 42], pedestrians [24], weapons [39, 40], or entirely metallic objects [30]. Through a new dataset and simulator, our work aims to expand the application of mmWave perception to a more open domain consisting of diverse categories.
@@ 3 Background
In this section, we provide a high-level overview on mmWave imaging and its properties.
@@
A mmWave radar transmits wireless signals with millimeter-wave wavelengths, captures their reflections, and uses them to generate 3D images, as shown in Fig. 3a. It uses two core signal processing techniques. First, it estimates the distance of each reflection by comparing the frequency between the transmitted and received signals. Second, it measures the angle of reflection (in both azimuth and elevation) by comparing the received signals between different antennas. 3D images can then be performed by calculating the reflection power emanating from each voxel in 3D space.
@@
where $I(x,y,z)$ is the complex mmWave image at location $(x,y,z)$, $K$ is the number of antennas, $N$ is the number of samples in each signal, $S_{j,k}$ is the jth sample of the received signal from the kth antenna, and $\lambda_{j}$ is the wavelength of the signal at sample j, which is proportional to the speed of light c and the frequency f as $\lambda=\frac{c}{f}$. $d(x,y,z,k)=2||(x,y,z)-p_{k}||$ is the round-trip distance from the kth antenna (at location $p_{k}$) to the point $(x,y,z)$.
3.1.1 Resolution
One important characteristic of a mmWave image is its resolution. The resolution in range (depth) is determined by the bandwidth, or the range of frequencies that the signal covers. We compute the depth resolution $\delta_{z}$ as:
@@ 3.1.1 Resolution
3.1.2 Synthetic Aperture Radar (SAR) Imaging
@@
3.2 Impact of Frequency
One important question is what frequency to use for imaging. The frequency introduces an important tradeoff. First, higher frequencies result in higher resolution (as seen in Eq. 3). On the other hand, higher frequency signals experience more attenuation, or power loss, as they travel through materials [5]. This means that high frequency signals may have higher noise through thick occlusions than lower frequencies. This is why our dataset and simulator include multi-spectral (multi-frequency) mmWave images.
When selecting frequencies, we chose mmWave signals because lower frequencies (e.g., WiFi) have poor resolution [9], while higher ones (e.g., Terahertz) struggle with occlusions [17]. In mmWave spectrum, we selected 24 GHz and 77 GHz since these are the bands approved by the FCC for industrial and commercial use [20] and most commercial mmWave radars operate within them.
3.3 Types of Reflections
mmWave signals primarily exhibit specular (mirror-like) reflections [32], as illustrated in Fig. 3. Due to the significantly longer wavelengths of mmWave signals compared to visible light (millimeters vs nanometers), surfaces that appear diffuse (i.e., omnidirectional) for visible light (blue) often appear specular for mmWave signals (red).
In addition to specular reflections, mmWaves also exhibit strong reflections when a wave strikes a sharp edge, scattering it in multiple directions. These reflections, known as edge diffraction, can impact signal propagation and perception differently from diffuse or specular reflections.
@@ 3.3 Types of Reflections
4 MITO Dataset
In this section, we describe the design choices and process of creating MITO. The dataset consists of over 24 million mmWave frames (where each frame is a complex-valued time series). We used these frames to generate 550 unique, high-resolution 3D mmWave SAR images of 76 diverse objects.
4.1 Object Selection
MITO builds on the YCB dataset [14], a standard dataset of everyday objects used for robotic manipulation. We select 76 out of 80 objects from the YCB dataset (the remaining 4 were discarded due to not being able to obtain the exact object). Unlike previous mmWave datasets, these objects cover a diverse range of shapes, sizes, materials, and categories. [^3] Additional details on object diversity are provided in supplementary. The YCB dataset also provides 2D images and 3D object meshes, which we use to simulate mmWave images.
4.2 Data Collection Protocol
4.2.1 Real-World mmWave Images
Fig. 1 visualizes our data collection setup. We connect a 77 GHz radar (TI IWR1443Boost [6] and DCA1000EVM [1]), 24 GHz radar (Infineon Position2Go [4]), and Intel Realsense D415 [25] to an end-effector of a UR5e robotic arm [44] using custom-designed 3D-printed parts. Since the robot and radars are controlled from different computers, we synchronize them by setting both computers to the same NTP time server.
We place the objects in two setups that minimize background reflections in different ways. In the first, we fix each object onto a tripod by hot-gluing a 3D-printed nut to each object and screwing it onto a threaded tripod. Because there is no surface directly below the object, any strong reflections from the background are avoided. In the second, we place the object directly on a styrofoam block, since styrofoam is largely transparent to mmWave signals [7]. We also include experiments where the tripod/styrofoam is placed in the scene without any objects so that the impact of the tripod/styrofoam can be removed by subtracting the empty image from the main image [10].
At the start of an experiment, the robot moves the camera to the middle of the workspace to take an RGB-D image prior to any potential occlusion (i.e., in the NLOS scenario). Then, it moves the radars throughout the workspace, covering a rectangular space. The aperture of our experiments range from 55cm by 22cm to 60cm by 45cm, ensuring that each experiment uses an aperture which covers the full object. The radars continuously measure mmWave signals while the robot is moving, and the robot records its trajectory over time. After we complete an experiment in LOS, we run an experiment in NLOS without moving the object. We place a layer of cardboard over the object to simulate the case where the object was placed inside a box. We also collect experiments with four layers of fabric over the cardboard to simulate more dense occlusions.
4.2.2 Synthetic mmWave Images
In addition to the real-world data, we also provide a simulation tool to generate synthetic mmWave images, which can be used to increase the volume of data to aid in model training. We include at least 2 different simulation images for each object, along with the simulator itself for producing additional images (see Sec. 5 for more details).
4.3 Data Processing
To process a real-world mmWave image, we need to know the location where each measurement was taken. To do this, we interpolate the timestamped robot locations to match the timestamp of the radar measurements. The interpolated locations, in addition to the raw radar measurements, are then processed into a 3D mmWave image by applying Eq. 1. Example images can be seen in Fig. 4, which shows the segmented RGB images (1st column), 77 GHz LOS (2nd column), 77 GHz NLOS (3rd column), 24 GHz LOS (4th column), and 24 GHz NLOS (5th column) mmWave images. Additional images are provided in the supplementary.
We produce 550 3D, complex-valued mmWave images at two different frequencies (24 GHz & 77 GHz) in both LOS and NLOS settings using 24 million raw measurements and locations. We include the raw data in MITO  to enable flexibility in leveraging and/or postprocessing the data using a custom pipeline. We also include images of objects at different angles and images with multiple objects.
For each mmWave image, we use its corresponding RGB-D image (in LOS) to provide a ground truth object segmentation mask. The segmentation masks are produced by using an interactive interface for the segment-anything model (SAM) [27] to select the points on the RGB picture as a prompt for the model. The human annotator continues selecting points until the mask covers the full object. We then use the camera’s location to transform and align the RGB-D ground-truth mask with the mmWave image. More details on the alignment process are in the supplementary.
4.4 Analysis of Different Frequencies
Recall from Sec. 3.2 that the frequency of the mmWave signal impacts both the resolution of the resulting image and the ability for the signal to traverse through occlusions. We analyze this phenomenon in real-world images in MITO in Fig. 5. Fig.5a-b show a 77 GHz and 24 GHz image of a padlock in LOS, respectively. The 77 GHz image is higher resolution, as expected. On the other hand, Fig. 5c-d show 77 GHz and 24 GHz images of the same object in NLOS (placed underneath cardboard). In this case, the 24 GHz image is very similar to the LOS case, while the 77 GHz shows some artifacts that were not present in LOS. This is because the 24 GHz signal travels through the cardboard relatively unaffected, while the 77 GHz is partially reflected and absorbed. We hope that by including two different frequencies, future algorithms can leverage both images to benefit from high-resolution and robustness to occlusions.
@@ 4.4 Analysis of Different Frequencies
5 MITO Simulator